OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
BACKBONE US
Incident Report for Network & Infrastructure
Resolved
We lost 2 fibers between ASH and BHS and 1 fiber between NWK and ASH.
Our optical team manage to bring back the fibers online, we are currently analyzing to find the root cause of this incident.


Update(s):

Date: 2018-05-24 09:24:59 UTC
The problem of loss was corrected by the provider this morning.

Date: 2018-05-23 04:47:51 UTC
The fiber provider will troubleshoot and fix the loss the 23 of May between 02h00 and 06h00 (Local time)

Date: 2018-05-21 12:07:08 UTC
There was an event between 2 amplifier sites on our darkfiber NWK<>ASH.
We lost 10dB on this span.

We are contacting the fiber provider to correct this loss.

Circuit are up and running.
Posted May 21, 2018 - 10:06 UTC